A MILLIMETER WAVE LINE SURVEY OF THE HIGH-MASS STAR-FORMING CORES NGC 6334 I AND I(NORTH)

\underline{SVEN THORWIRTH}, I. Physikalisches Institut, Universität zu Köln, Zülpicher Str. 77, 50937 Köln, Germany; ACHIM R. TIEFTRUNK, Deutsche Forschungsgemeinschaft, Kennedyallee 40, 53175 Bonn, Germany.

A millimeter line survey has been conducted towards the compact high mass star forming cores NGC 6334 I and I(North) using the 15 m Swedish ESO Submillimeter Telescope, SEST, located on La Silla, Chile^{a}. Especially NGC 6334 I shows a plethora of molecular lines, indicative of a highly active chemistry. The advantage over classical sources such as Orion and SgrB2 is the narrow line profiles towards these cores. Thus, more lines can be detected before the confusion limit is reached making NGC 6334 I a prime target in future searches of “new” interstellar molecules.

aS. Thorwirth, G. Winnewisser, S. T. Megeath, and A. R. Tieftrunk, “The Twin-Core System NGC6334 I & I(N): High-Mass Stars and Stellar Clusters in the Making”, in Galactic Star Formation Across the Stellar Mass Spectrum, eds. J. De Buizer & N.S.

van der Bliek, ASP Conference Series 287, 257 (2003).